I am working with subtraction of real numbers. I can not seem to grasp this. All my work on my own has been wrong and I do not understand how to do these.
Answer by richard1234(7193)

It's similar to the subtraction you learned in elementary school, e.g. 10-6 = 4. However it's more general since you can allow for negative numbers, fractions/decimals, etc. as in the following examples:

5 - 8 = -3
5 - (-8) = 5+8
-3.8 - (-2) = -3.8 + 2 = -1.8
-104 - 1 = -105
1/2 - 1/4 = 2/4 - 1/4 = 1/4
2/5 - (-3/10) = 2/5 + 3/10 = 4/10 + 3/10 = 7/10
1/3 - 4/9 = 3/9 - 4/9 = -1/9